### v1

# A Fractional Calculus Framework for Open Quantum Dynamics: From Liouville to Lindblad to Memory Kernels

Abstract
Open quantum systems exhibit dynamics ranging from purely unitary evolution to irreversible dissipative relaxation. The Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation uniquely characterizes Markovian dynamics that are completely positive and trace-preserving (CPTP), yet many physical systems display non-Markovian features such as algebraic relaxation and coherence backflow beyond the reach of semigroup evolution. Fractional calculus provides a natural framework for describing such long-memory behavior through power-law temporal kernels introduced by fractional time derivatives. Here we establish a unified hierarchy that embeds fractional quantum master equations within the broader landscape of open system dynamics. The fractional master equation forms a structured subclass of memory-kernel models, reducing to the GKSL form at unit fractional order. Through Bochner–Phillips subordination, fractional evolution is expressed as an average over Lindblad semigroups weighted by a power-law waiting-time distribution. This construction ensures physical consistency, explains the algebraic origin of long-time decay, and bridges unitary, Markovian, and structured non-Markovian regimes. The resulting framework positions fractional calculus as a rigorous and unifying language for quantum dynamics with intrinsic memory, enabling new directions for theoretical analysis and quantum simulation.
preprint: AIP/123-QED
I Introduction
The dynamics of an isolated quantum system are governed by the Liouville–von Neumann equation, which generates perfectly unitary and reversible evolution vonNeumann1929. When a system interacts with its environment, however, the reduced dynamics generally become irreversible and often highly nontrivial. The celebrated Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) equation Gorini1976; Lindblad1976 provides the unique characterization of Markovian open system dynamics, ensuring CPTP semigroup evolution Alicki1977. Yet, many experimental platforms–from superconducting qubits Murch2013; Kjaergaard2020, nitrogen-vacancy centers PhysRevLett.121.060401, trapped ions Clos2016, and ultracold gases PhysRevA.87.012127—display non-Markovian behavior such as non-exponential decay, coherence backflow, and power-law tails, indicating the presence of memory effects beyond the reach of standard Lindblad dynamics Vacchini2011; Dominy2015; Chruscinski2017.
Fractional calculus has emerged as a powerful framework for modeling such long-memory and anomalous relaxation phenomena Hilfer2000; Tarasov2021. By replacing the first-order derivative with a fractional time derivative, one introduces a power-law memory kernel that endows the dynamics with nonlocal temporal correlations. The corresponding decay, governed by the Mittag–Leffler function Mainardi2000, interpolates continuously between exponential and algebraic regimes. Despite its empirical success in describing slow relaxation and heavy-tailed dynamics, most fractional formulations in quantum contexts remain phenomenological, lacking a rigorous connection to established open system theory Breuer2002. Fundamental questions persist regarding how fractional evolution relates to the GKSL semigroup, whether it preserves complete positivity, and how it fits within the general class of memory-kernel quantum master equations (QMEs).
In this work, we develop a unified theoretical framework that places fractional dynamics within the broader hierarchy of quantum open system formalisms. We show that fractional master equations form a structured subclass of non-Markovian QMEs, coinciding with standard Lindblad dynamics in the limit of unit fractional order. This perspective reveals that fractional differentiation introduces memory at the level of the generator rather than through explicit environmental coupling, bridging the conceptual gap between purely unitary and dissipative descriptions.
Together, these insights establish a coherent storyline that connects the principal regimes of quantum dynamics: unitary Liouville evolution forms the base, Lindbladian semigroups represent Markovian open systems, and fractional extensions introduce structured non-Markovianity through power-law memory kernels. This unified framework provides both conceptual clarity and a mathematical foundation for simulating long-memory quantum processes on emerging quantum and hybrid computing platforms. Unlike phenomenological fractional models, the present formulation embeds fractional differentiation directly within the GKSL generator and preserves complete positivity through Bochner–Phillips subordination Bochner1955; Feller1971. This distinguishes it from projection-operator Nakajima1958; Zwanzig1960 and time-convolutionless approaches Breuer2002; Breuer2016, which introduce memory through explicit environmental couplings rather than at the generator level.
II Theory
II.1 Preliminaries
Let $\mathcal{H}$ be a separable Hilbert space, and let $\mathcal{B}_{1}(\mathcal{H})$ denote the Banach space of trace-class operators acting on $\mathcal{H}$. A physical quantum state is a density operator $\rho\in\mathcal{B}_{1}(\mathcal{H})$ satisfying $\rho\geq 0$ and $\mathrm{Tr}(\rho)=1$. The time evolution of $\rho$ is represented by a family of linear maps $\{\Phi(t)\}_{t\geq 0}$ acting on $\mathcal{B}_{1}(\mathcal{H})$. Physical consistency requires that each $\Phi(t)$ be CPTP; this condition ensures that the evolution of $\rho$ remains valid even when the system is part of a larger entangled state, and is thus central to any legitimate description of quantum dynamics Nielsen2000; Chruscinski2017.
When the system is isolated, the evolution is unitary and memoryless, governed by the Liouville–von Neumann equation vonNeumann1929. The corresponding maps form a one-parameter unitary group that is CPTP, invertible, and norm-preserving. In the presence of environmental coupling, irreversibility and decoherence arise. If the dynamics remain time-homogeneous and memoryless, the family $\{\Phi(t)\}$ constitutes a quantum dynamical semigroup Alicki1977:
$\Phi(0)=\mathbb{I},\qquad\Phi(t+s)=\Phi(t)\Phi(s),$ (1)
with $\Phi(t)$ CPTP and norm-continuous. The Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) theorem guarantees that any such semigroup admits a generator $\mathcal{L}$ of Lindblad form Gorini1976; Lindblad1976, uniquely characterizing Markovian open system dynamics in which all memory of past states is lost.
In many physical systems, this Markovian approximation breaks down. Coherence revivals, stretched-exponential relaxation, and algebraic long-time tails all indicate that the generator $\mathcal{L}$ becomes effectively time-nonlocal Breuer2016; Li2018; deVega2017. The resulting dynamics are better described by memory-kernel QMEs, which introduce explicit temporal correlations through convolution integrals Nakajima1958; Zwanzig1960. Fractional calculus provides a compact and rigorous formulation of such dynamics: fractional time derivatives generate power-law kernels that naturally give rise to Mittag–Leffler relaxation, interpolating between exponential and algebraic behavior Mainardi2000; Gorenflo2014; Luchko2020. In the limit $\alpha~\to~1$, the fractional derivative recovers the Markovian Lindblad semigroup; for $\alpha<1$, the dynamics become non-divisible yet remain CPTP and physically consistent Chruscinski2017; Hall2014.
Fractional dynamics therefore provide a unified mathematical bridge connecting distinct regimes of quantum evolution—from unitary to non-unitary, from memoryless to memoryful—under a single operator-theoretic framework. To make this connection precise, we introduce the following hierarchy of dynamical models.
Definition 1 (Liouville Dynamics).
Liouville dynamics describe the unitary evolution of a closed quantum system governed by a Hermitian Hamiltonian $H=H^{\dagger}$:
$\dot{\rho}(t)=-i[H,\rho(t)].$ (2)
The maps $\Phi(t)(\rho)=e^{-iHt}\rho e^{iHt}$ form a one-parameter group of CPTP and norm-preserving transformations.
Definition 2 (Lindblad/GKSL Dynamics).
Lindblad dynamics generalize Liouville dynamics to Markovian open systems. They satisfy
$\dot{\rho}(t)=\mathcal{L}\rho(t),$ (3)
with
$\mathcal{L}(\rho)=-i[H,\rho]+\sum_{j}\gamma_{j}~\left(L_{j}\rho L_{j}^{\dagger}-\tfrac{1}{2}\{L_{j}^{\dagger}L_{j},\rho\}\right).$ (4)
Here, $\gamma_{j}\geq 0$ quantifies the strength of the dissipative channel represented by the jump operator $L_{j}$. The solution $\rho(t)=e^{t\mathcal{L}}\rho(0)$ defines a CPTP semigroup, and complete positivity is ensured precisely by the Lindblad structure of $\mathcal{L}$ superoperator.
Definition 3 (Memory-Kernel Quantum Master Equation).
A convolution-type or memory-kernel quantum master equation describes non-Markovian dynamics through an integral kernel $K(t-\tau)$:
$\dot{\rho}(t)=\int_{0}^{t}K(t-\tau)~\mathcal{L}\rho(\tau)~d\tau.$ (5)
The kernel encodes temporal correlations and environmental backaction. Choosing $K(t)=\delta(t)$ recovers Lindblad dynamics, while more general kernels yield stretched-exponential or power-law relaxation Breuer2002; deVega2017.
Definition 4 (Fractional Master Equation).
For $0<\alpha<1$, a fractional master equation replaces the first-order time derivative with a fractional (Caputo) derivative of order $\alpha$:
$\prescript{C}{}{D}_{t}^{\alpha}\rho(t)=\mathcal{L}\rho(t).$ (6)
Here $\mathcal{L}$ is a densely defined linear generator on $\mathcal{B}_{1}(\mathcal{H})$. The fractional derivative introduces a long-range power-law memory kernel, giving rise to Mittag–Leffler relaxation that interpolates between exponential and algebraic decay. In the limit $\alpha~\to~1$, the equation reduces to the standard first-order master equation $\dot{\rho}(t)=\mathcal{L}\rho(t)$ Podlubny1999; Tarasov2021.
Remark 1 (Fractional Liouville and Fractional Lindblad).
(i) Fractional Liouville dynamics correspond to the closed-system case where $\mathcal{L}=-i[H,~\cdot~]$ with $H=H^{\dagger}$. They describe coherent yet memoryful evolution that remains CPTP but non-divisible, effectively a continuous mixture of unitaries Bochner1955; Feller1971.
(ii) Fractional Lindblad dynamics arise when $\mathcal{L}$ takes the GKSL form. These dynamics represent physically consistent non-Markovian extensions of standard Lindbladian evolution, equivalent to subordinated Lindblad semigroups governed by power-law waiting-time distributions Hall2014.
Liouville $\alpha=1,\ \lambda=0$ Unitary, reversible Markovian Lindblad $\alpha=1,\ \lambda>0$ CPTP semigroup Non-Markovian QME $\alpha\!\neq\!1,\ \lambda>0$ Not always CPTP Fractional Lindblad $\prescript{C}{}{D}_{t}^{\alpha}\rho=\lambda\,\mathcal{L}\rho$ $0<\alpha<1,\ \lambda>0$ CPTP, power-law memory $\lambda\downarrow,~\alpha\uparrow$$\alpha\uparrowgeneral$ memory
Figure 1: Schematic hierarchy of quantum dynamical regimes unified by fractional subordination. The parameters $\alpha$ (memory order) and $\lambda$ (dissipation strength) interpolate smoothly between unitary Liouville dynamics ($\alpha{=}1,\lambda{=}0$), Markovian Lindblad semigroups ($\alpha{=}1,\lambda{>}0$), and general memory–kernel quantum master equations ($\alpha{\neq}1,\lambda{>}0$). Fractional or subordinated Lindblad dynamics ($0{<}\alpha{<}1,\lambda{>}0$) form the CPTP bridge connecting these regimes.
These nested definitions establish the structural hierarchy as schematically illustrated in Figure 1, in which fractional calculus connects the memoryless semigroup of Lindblad evolution to the nonlocal, algebraic relaxation characteristic of non-Markovian quantum systems Breuer2016; Li2018; deVega2017.
The structural hierarchy above admits a natural mathematical interpretation through the concept of subordination. In this picture, fractional dynamics arise as averages over standard semigroup evolutions with respect to a probability density in operational time Bochner1955; Feller1971. Specifically, the fractional flow can be represented as a Bochner–Phillips subordination of a Lindblad semigroup, in which the physical time $t$ is replaced by a random operational time distributed according to a power-law waiting-time function Bochner1955; Feller1971; Mainardi2000. This representation guarantees complete positivity whenever the underlying generator $\mathcal{L}$ is of GKSL form and provides a direct link between fractional differentiation, non-divisibility, and long-memory behavior Rivas2014. In the next section, we make this correspondence explicit and examine how fractional subordination generates algebraic relaxation and interpolates smoothly between unitary, Markovian, and non-Markovian regimes.
Fractional extensions of quantum master equations have been explored in various contexts, from phenomenological models of quantum dissipation with power-law memory Tarasov2021 to more recent formulations of time-fractional Lindblad dynamics for weakly coupled systems Ang2024MemoryInducedWeakDissipation. Parallel developments have applied subordination principles in quantum settings, including Bochner and topological subordination for fractional evolution equations Feller1971; Sandev2018FCAA. These approaches, while illuminating, often treat fractional derivatives or memory kernels as postulated modifications rather than consequences of a well-defined stochastic process. In contrast, the present work provides a rigorous unification based on Bochner–Phillips subordination of Lindblad semigroups. This construction ensures that the fractional propagator CPTP for any $0<\alpha\leq 1$, thereby reconciling fractional dynamics with quantum dynamical semigroup theory. The probabilistic interpretation in terms of random operational times $u\sim f_{\alpha}(u,t)$ further connects power-law relaxation directly to waiting-time statistics, clarifying the origin of algebraic tails in open system coherence.
III Fractional Master Equation
Fractional calculus provides a rigorous way to encode long-range memory in quantum dynamics. In the fractional master equation (6), $\mathcal{L}$ is a densely defined generator on $\mathcal{B}_{1}(\mathcal{H})$, and the Caputo derivative acts on differentiable functions $f:[0,\infty)\to\mathcal{B}_{1}(\mathcal{H})$ as
$\prescript{C}{}{D}_{t}^{\alpha}f(t)$ $=\Big(I^{1-\alpha}\circ\frac{d}{dt}\Big)f(t)$
$=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}\frac{f^{\prime}(\tau)}{(t-\tau)^{\alpha}}~d\tau,\qquad 0<\alpha<1,$
$\equiv\int\kappa_{C}(t-\tau)f^{\prime}(\tau)d\tau,$ (7)
where $\kappa_{C}(t)=\frac{t^{-\alpha}}{\Gamma(1-\alpha)}$ and $I^{1-\alpha}$ denotes fractional integration Podlubny1999; Kilbas2006; Luchko2020. The Gamma function
[…]
fixes the normalization of the weakly singular kernel $(t-\tau)^{-\alpha}$ in accordance with the Riemann–Liouville (RL) conventions Kilbas2006. In this way, an explicit power-law weighting of past rates $f^{\prime}(\tau)$ replaces instantaneous evolution by a history-dependent response.
III.1 Equivalence to convolution master equations
Starting from Eq. (6), applying the RL integral $I_{t}^{\alpha}$ and using $\prescript{C}{}{D}_{t}^{\alpha}I_{t}^{\alpha}=\mathbb{I}$ gives the Volterra form
$\rho(t)=\rho(0)+\int_{0}^{t}K_{\alpha}^{(\mathrm{V})}(t-\tau)~\mathcal{L}\rho(\tau)~d\tau$ (9)
with the kernel
$K_{\alpha}^{(\mathrm{V})}(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}.$ (10)
Differentiating Eq. (9) (Leibniz rule for weakly singular kernels) yields the differential convolution form
$\dot{\rho}(t)=\int_{0}^{t}k_{\alpha}(t-\tau)~\mathcal{L}\rho(\tau)~d\tau+\frac{t^{\alpha-1}}{\Gamma(\alpha)}~\mathcal{L}\rho(0)$ (11)
with the kernel
$k_{\alpha}(t)=\frac{d}{dt}K_{\alpha}^{(\mathrm{V})}(t)=\frac{t^{\alpha-2}}{\Gamma(\alpha-1)}.$ (12)
If $\mathcal{L}\rho(0)=0$ (for instance, when $\rho(0)$ is a fixed point of $\mathcal{L}$), the boundary term vanishes and Eq. (11) reduces to a pure convolution master equation Chruscinski2017; Tarasov2021.
It is helpful to separate the roles of the three kernels that appear across these equivalent forms. The Volterra kernel $K_{\alpha}^{(\mathrm{V})}(t)=t^{\alpha-1}/\Gamma(\alpha)$ multiplies $\mathcal{L}\rho$ in the integral equation Eq. (9) and is completely monotone on $\mathbb{R}_{+}$; this complete monotonicity means $K_{\alpha}^{(\mathrm{V})}$ is the Laplace transform of a positive measure and underlies the subordination representation developed later. Differentiating $K_{\alpha}^{(\mathrm{V})}$ produces $k_{\alpha}(t)=t^{\alpha-2}/\Gamma(\alpha-1)$, the kernel in the differential convolution form Eq. (11). By contrast, the Caputo derivative itself can be written as a convolution of $\dot{\rho}$ with $\kappa_{C}(t)=t^{-\alpha}/\Gamma(1-\alpha)$; this kernel lives inside the Caputo operator and should not be identified with the memory kernels that multiply $\mathcal{L}\rho$. Detailed proofs are provided in Appendix A (Proposition 1).
III.2 Mittag–Leffler solution and physical interpretation
The formal solution of the fractional master equation can be written using the operator Mittag–Leffler function as
$\rho(t)=E_{\alpha}(t^{\alpha}\mathcal{L})\,\rho(0),$ (13)
where the Mittag-Leffler function $E_{\alpha}(z)=\sum_{n=0}^{\infty}z^{n}/\Gamma(1+\alpha n)$ is defined via its convergent power series Mainardi2000. If $\mathcal{L}$ is diagonalizable with eigenvalues $\{\lambda_{j}\}$ and spectral projectors $\{\Pi_{j}\}$, this reduces to
$\rho(t)=\sum_{j}E_{\alpha}(\lambda_{j}t^{\alpha})\,\Pi_{j}\rho(0).$ (14)
For non-diagonalizable $\mathcal{L}$, the same operator expression (13) remains valid and the action on generalized eigenspaces simply involves polynomially weighted derivatives of $E_{\alpha}$ (Jordan–block structure). Notably, the subordination representation and complete monotonicity of $K_{\alpha}^{(\mathrm{V})}$ do not require diagonalizability. For $\alpha=1$, one recovers $E_{1}(z)=e^{z}$ and the familiar exponential Lindblad relaxation. For $0<\alpha<1$, the response crosses over from stretched-exponential at short times to algebraic at long times, capturing persistent memory and non-Markovian relaxation observed in many physical settings Metzler2014.
Because $K_{\alpha}^{(\mathrm{V})}$ is completely monotone, the dynamical map remains CPTP whenever $\mathcal{L}$ is of GKSL form; however, the resulting family generally lacks the semigroup property and is not CP-divisible Rivas2014. The next subsection makes this precise via the Bochner–Phillips subordination formula, in which physical time is replaced by a random operational time with power-law waiting statistics. Additional details on the Caputo–Volterra equivalence and on fractional Adams–Moulton discretizations appear in Appendices A and B. Recent analyses of driven two-level systems show analogous fractional relaxation and coherence damping governed by Mittag–Leffler functions, extending the Rabi problem to fractional time evolution lopez2025generalisedfractionalrabiproblem.
III.3 Fractional Subordination
Fractional dynamics is equivalent to running the usual Markovian evolution $e^{u\mathcal{L}}$ on a random clock $u$ whose statistics encode memory; averaging over that clock yields the fractional map.
$\hat{\rho}(s)=s^{\alpha-1}\big(s^{\alpha}\mathbb{I}-\mathcal{L}\big)^{-1}\rho(0).$ (15)
Using Bochner’s functional calculus for the GKSL generator $\mathcal{L}$ Bochner1955; Feller1971; BaeumerMeerschaert2001; Schilling2012Bernstein; MeerschaertSikorskii2012,
$\big(s^{\alpha}\mathbb{I}-\mathcal{L}\big)^{-1}=\int_{0}^{\infty}e^{-us^{\alpha}}~e^{u\mathcal{L}}~du,$ (16)
and inverting the Laplace transform introduces the inverse-stable (Lévy) density $f_{\alpha}(u,t)$:
$\rho(t)=\int_{0}^{\infty}f_{\alpha}(u,t)~e^{u\mathcal{L}}\rho(0)~du.$ (17)
The kernel satisfies $f_{\alpha}(u,t)\geq 0$ and $\int_{0}^{\infty}f_{\alpha}(u,t)~du=1$, which makes the evolution (17) represents a convex mixture of CPTP semigroups $e^{u\mathcal{L}}$ (detailed in Appendix C). Specifically, the subordination integral in Eq. (17) represents a probabilistic construction of fractional dynamics: the inverse–stable density $f_{\alpha}(u,t)\!\geq 0$ acts as a genuine probability distribution for the random operational time $u$ of an underlying stochastic clock. Consequently, $\Phi_{\alpha}(t)$ can be interpreted as the statistical mixture $\Phi_{\alpha}(t)=\mathbb{E}_{U(t)}[e^{U(t)\mathcal{L}}]$, i.e., an ensemble average of standard Markovian (Lindblad) evolutions sampled at random times $U(t)$ with power–law waiting statistics. This probabilistic representation ensures that complete positivity is preserved and explains how algebraic relaxation emerges naturally from the statistics of random operational time.
The representation (17) also has an important algorithmic interpretation. Because $e^{u\mathcal{L}}$ admits standard quantum–trajectory and quantum–jump unravelings, the subordination formula shows that any Markovian trajectory method can be extended to fractional, non-Markovian dynamics simply by replacing the physical time by a random clock $U(t)$ drawn from $f_{\alpha}(u,t)$. Thus, fractional dynamics corresponds to sampling ordinary Lindbladian trajectories at random operational times, with the memory effects arising entirely from the statistics of $U(t)$.
In the special case $\mathcal{L}=-i[H,~\cdot~]$, Eq. (17) reduces to a fractional Liouville evolution,
$\rho(t)=\!\int_{0}^{\infty}\!f_{\alpha}(u,t)\,e^{-iHu}\rho(0)e^{iHu}\,du,$ (18)
which represents a convex mixture of unitary maps parameterized by the random operational time $u$. Because $f_{\alpha}(u,t)$ is non-negative and normalized, this evolution is CPTP for all $\alpha$, continuously connecting the coherent, memoryful fractional regime to the strictly unitary Liouville dynamics recovered in the limit $\alpha\!\to\!1$. This establishes the Liouville equation as the coherent base of the hierarchy spanning unitary, Markovian, and non-Markovian quantum evolution.
Note that, one can plug Eq. (16) into Eq. (15), where the obtained Laplace kernel $s^{\alpha-1}e^{-us^{\alpha}}$ is completely monotone, implying that $f_{\alpha}(u,t)$ is a probability density on $u\geq 0$. Consequently, the subordinated map
$\Phi_{\alpha}(t):\ \rho(0)\mapsto\rho(t)$ (19)
preserves complete positivity and trace, i.e., remains CPTP whenever the underlying generator $\mathcal{L}$ is of GKSL form Bochner1955; Feller1971; Hall2014; Chruscinski2017. In standard (Markovian) open dynamics, the family $\{\Phi_{1}(t)\}_{t\geq 0}=\{e^{t\mathcal{L}}\}$ forms a semigroup obeying the composition law, Eq. (1); equivalently, for any $0\leq\tau\leq t$ there exists a CPTP map $\Lambda(t,\tau)$ such that
$\Phi_{1}(t)=\Lambda(t,\tau)~\Phi_{1}(\tau)\qquad\text{(CP-divisibility)}.$ (20)
By contrast, the fractional dynamics ($0<\alpha<1$) averages over random clock times in Eq. (17) and thereby breaks exact time composition in general:
$\Phi_{\alpha}(t)\neq\Phi_{\alpha}(t-\tau)~\Phi_{\alpha}(\tau),\qquad 0\leq\tau<t,$ (21)
so $\{\Phi_{\alpha}(t)\}$ is typically not a semigroup. A simple test follows from the known action on eigenoperators of $\mathcal{L}$:
$\mathcal{L}X=\lambda X\ \Rightarrow\ \Phi_{\alpha}(t)X=E_{\alpha}(\lambda t^{\alpha})~X.$ (22)
If $\{\Phi_{\alpha}(t)\}$ were a semigroup, one would require $E_{\alpha}(\lambda t^{\alpha})=e^{\mu t}$ for all $t$, which is impossible unless $\alpha=1$ (or $\lambda=0$). Hence, for $0<\alpha<1$ and nontrivial $\mathcal{L}$, the fractional family is not a semigroup and is generally not CP-divisible. In special commuting cases (e.g. pure dephasing channels) where all maps diagonalize simultaneously, intermediate CPTP maps may still exist, see, for example, Refs. PhysRevLett.112.120404; Sandev_2019; Ang2024MemoryInducedWeakDissipation).
For $\lambda<0$, the Mittag–Leffler factor obeys $E_{\alpha}(\lambda t^{\alpha})\sim-\big[\lambda~\Gamma(1-\alpha)\big]^{-1}t^{-\alpha}$ as $t\to\infty$. Populations and coherences therefore relax as a power law $t^{-\alpha}$ rather than exponentially, an experimentally relevant signature of long memory Weiss2012; Mainardi2000. In short, subordination guarantees CPTP at each time $t$ while replacing the Markovian semigroup structure with a memory-induced, non-composable flow that exhibits algebraic relaxation.
IV Numerical Demonstration and Discussion
We work in the exactly solvable pure–dephasing limit of the spin–boson model,
$H=\frac{\epsilon}{2}~\sigma_{z}+\sum_{k}\omega_{k}~b_{k}^{\dagger}b_{k}+\sigma_{z}\sum_{k}g_{k}~(b_{k}^{\dagger}+b_{k}),$ (23)
with independent boson modes $\{b_{k}\}$ in thermal equilibrium at inverse temperature $\beta$. Since $H,\sigma_{z}]=0$, populations are constants of motion; only the off–diagonal element (coherence) decays. If we assume an initially fractorized state for the total spin-boson system
$\rho_{\text{tot}}(0)=\rho_{S}(0)\otimes\rho_{B},$ (24)
with bath state $\rho_{B}=\frac{e^{-\beta H_{B}}}{Z_{B}}$ and bath Hamiltonian $H_{B}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$. Since $\rho_{B}$ is exactly a Gaussian state, tracing over a thermal Gaussian bath yields the exact coherence (Kubo–Martin–Schwinger identity) Leggett1987; Weiss2012; deVega2017,
$u(t)=\langle\sigma_{+}(t)\rangle=\langle\sigma_{+}(0)\rangle~e^{-i\epsilon t}~e^{-Q(t)},$ (25)
where $\sigma_{+}=\tfrac{1}{2}(\sigma_{x}+i\sigma_{y})$ and
[…]
with $J(\omega)$ the spectral density. A standard choice of $J(\omega)$ is the algebraic (sub–Ohmic/Ohmic/super–Ohmic) family with exponential cutoff Breuer2002; Weiss2012,
$J(\omega)=\eta~\omega^{\chi}\omega_{c}^{~1-\chi}e^{-\omega/\omega_{c}},\quad\chi\in\mathbb{R},~\eta>0,~\omega_{c}>0.$ (27)
as used in standard treatments of dissipative two-level systems and spin–boson dynamics Grifoni1998; Weiss2012. Table 1 provides the asymptotic scaling for the coherence that smoothly interpolates between Gaussian short-time decay, stretched-exponential-like intermediate behavior, and algebraic or saturating long-time tails, depending on $s$ and $t$ Leggett1987; Grifoni1998; Weiss2012; deVega2017. The real-time propagation of the dephasing functional $Q(t)$ and corresponding coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for $t\leq 0.2/\omega_{c}$ and $t\geq 5/\omega_{c}$ and three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are shown in Figure 2. In the weak-coupling and pure-dephasing limit, the reduced coherence $u(t)$ obeys a Nakajima–Zwanzig (NZ) convolution equation Nakajima1958; Zwanzig1960,
$\dot{u}_{\mathrm{NZ}}(t)=-\int_{0}^{t}\kappa_{\mathrm{NZ}}(t-\tau)u_{\mathrm{NZ}}(\tau)d\tau,$ (28)
where the memory kernel $\kappa_{\mathrm{NZ}}(t)$ is determined by the bath correlation function
$\kappa_{\rm NZ}\propto C(t)$ $=\langle B(t)B(0)\rangle$
[…]
with
$B(t)=\sum_{k}g_{k}\big(b_{k}^{\dagger}e^{i\omega_{k}t}+b_{k}e^{-i\omega_{k}t}\big),$ (30)
corresponding to the bath coupling operator. The exact solution,
$u_{\mathrm{NZ}}(t)=e^{-Q(t)}.$ (31)
coincides with the Gaussian-bath result in Eq. (25), confirming that $Q(t)$ fully characterizes the non-Markovian dephasing kernel.
Table 1: Short- and long-time asymptotic forms of the dephasing functional $Q(t)$ and corresponding coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for different bath exponents $s$. Here, the asymptotic forms in each column follow standard analyses of the spin–boson model, and focus on the non-thermal quantum decoherence, thus $\coth(\beta\omega/2)\sim 1$. For $|\langle\sigma_{+}(t)\rangle|$ in short-time Gaussian decay, $t_{G}=\sqrt{2}~\omega_{c}^{-1}(\eta\,\Gamma(\chi+1))^{-1/2}$. For $|\langle\sigma_{+}(t)\rangle|$ in long-time sub-ohmic case, $C_{\chi}(\omega_{c})=\frac{2}{\pi}\,\eta\,\omega_{c}^{\,1-\chi}\Gamma(1-\chi)\sin\!\Big(\frac{\pi\chi}{2}\Big)$. For $|\langle\sigma_{+}(t)\rangle|$ in the long-time super-Ohmic case, $Q_{\infty}=\tfrac{2}{\pi}\eta\,\Gamma(\chi\!-\!1)$.
Short-time Leggett1987; Loss1998; Weiss2012 Long-time ($t\gg 1/\omega_{c}$)
($t\ll 1/\omega_{c}$) Sub–Ohmic ($\chi<1$Leggett1987; Grifoni1998 Ohmic ($\chi=1$Leggett1987; Weiss2012; deVega2017 Super–Ohmic ($\chi>1$Grifoni1998; Weiss2012
$Q(t)$ $\mathcal{O}\left(\frac{1}{2}\eta\,\Gamma(\chi+1)\,\omega_{c}^{2}\,t^{2}\right)$ $\mathcal{O}\left(C_{\chi}(\omega_{c})\,t^{1-\chi}\right)$ $\mathcal{O}\left(\frac{\eta}{2}\ln(\omega_{c}^{2}t^{2})\right)$ $\mathcal{O}\left(Q_{\infty}-D_{\chi}\,t^{-(\chi-1)}\right)$
$|\langle\sigma_{+}(t)\rangle|$ $\mathcal{O}\left(e^{-t^{2}/t_{G}^{2}}\right)$ $\mathcal{O}\left(e^{-\omega_{c}^{1-\chi}\,t^{1-\chi}}\right)$ $\mathcal{O}\left(t^{-\eta}\right)$ $\mathcal{O}\left(e^{-Q_{\infty}}\!\big[1+D_{\chi}\,t^{-(\chi-1)}\big]\right)$
Figure 2: Short-time ($t\leq 0.2/\omega_{c}$, left panels) and long-time ($t\geq 5/\omega_{c}$, right panels) asymptotic propagation of the dephasing functional $Q(t)$ and corresponding coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for three different bath exponents $\chi=0.5$, $1.0$, and $1.5$. Solid lines denote numerical integrations of Eq. (26); open circles show asymptotic expressions (Table. 1) with optimized prefactors, consistent with the asymptotic analysis in Refs. Leggett1987; Grifoni1998; Weiss2012; deVega2017.
IV.1 Why Markovian dephasing misses $e^{-Q(t)}$?
A common Markovian dephasing of the spin-boson model relies on the time‑homogeneous Lindbladian model (3) and (4) with $\gamma_{j}\geq 0$. Here, the system Hamiltonian $H_{S}=\frac{\epsilon}{2}\sigma_{z}$ and a single jump operator $L=\sigma_{z}$ is employed. This Markovian model can then be transformed to
$\dot{u}_{\rm M}(t)=(i\epsilon-2\gamma)\,u_{\rm M}(t)$ (32)
that yields an exponential dephasing
$u_{\rm M}(t)=u_{\rm M}(0)\,e^{i\epsilon t}\,e^{-2\gamma t}.$ (33)
Eq. (32) can also be written in the NZ form
$\dot{u}_{\rm M}(t)=-\int_{0}^{t}\kappa_{\rm M}(t-\tau)\,u_{\rm M}(\tau)\,d\tau$ (34)
with a $\delta$-kernel,
$\kappa_{\rm M}(t)=2\gamma\delta(t).$ (35)
Since the entire Markovian kernel is a $\delta$-function concentrated at the present time, this is different from the exact kernel (29) that possesses a long tail inherited from the bath spectrum $J(\omega)$. Therefore, the time-homogeneous Lindbladian cannot capture the long-time memory that produces $u(t)=e^{-Q(t)}$. The “missing terms” are precisely the nonlocal (finite-memory) contributions. It is worth noting that, in order to reproduce $u(t)=e^{-Q(t)}$, one can choose the rate $\gamma$ to be time-dependent. For example, with $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$, we have
$\dot{u}(t)=\big(i\epsilon-2\gamma(t)\big)\,u(t)~~\Rightarrow~~u(t)=u(0)\,e^{i\epsilon t}\,\exp\!\Big(-2\!\int_{0}^{t}\!\gamma(\tau)\,d\tau\Big).$ (36)
This more flexible time-local dephasing model merely re‑encodes the exact memory in $\gamma(t)$ through $Q(t)$ to reproduce the exact dephasing law. Moreover, if $\gamma(t)$ were to become temporarily negative (possible for structured environments), the dynamics is non-Markovian (not CP-divisible) Rivas2014; Hall2014, and the GKSL form with constant positive rates is no longer valid.
Figure 3: Comparison of exact coherence from Figure 2 with Markovian and non-Markovian fits. The asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for three different bath exponents $\chi=0.5$, $1.0$, and $1.5$ are computed by constant $\gamma$ (Markovian, left panel) and time-local $\gamma(t)=\tfrac{1}{2}\dot{Q}(t)$ (non-Markovian, right panel), respectively. The constant rate $\gamma$ is obtained by fitting $\log|\langle\sigma_{+}(t)\rangle|$ to a linear decay in the intermediate-time regime $t\in[2/\omega_{c},60/\omega_{c}]$ (blue shade, left panel), while solid lines are obtained by numerically integrating Eq. (26).
The analysis above highlights that a constant-rate Lindbladian cannot capture the long-tailed memory encoded in $Q(t)$, while a time-dependent rate merely re-parametrizes this memory without providing structural insight. A more fundamental generalization is therefore to embed the memory directly into the generator through fractional differentiation. We next show that replacing the first-order derivative in the Lindblad equation with a fractional derivative produces a compact, CPTP, and physically transparent model that reproduces the exact non-Markovian coherence $e^{-Q(t)}$ across spectral regimes.
Figure 4: Asymptotic propagation of the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ for the spin–boson model obtained from the single–order fractional propagation and compared with the numerically exact result (solid lines, computed by direct integration of Eq. (26)). Results are shown for three spectral exponents, $\chi=0.5$, $1.0$, and $1.5$. The fractional parameters $(\alpha,\lambda)$ are determined by the least–squares fitting to the exact propagation within the indicated intermediate–time windows (blue shades, left panels): $t\!\in\![2/\omega_{c},20/\omega_{c}]$ (top), $t\!\in\![2/\omega_{c},40/\omega_{c}]$ (middle), and $t\!\in\![2/\omega_{c},60/\omega_{c}]$ (bottom). Circle markers denote the fractional results; the shaded regions mark the fitting intervals.
IV.2 Single–order fractional Lindbladian for structured baths
The failure of the conventional Markovian Lindbladian to reproduce the nonexponential coherence $e^{-Q(t)}$ underscores the need for a minimal generalization that can capture the memory effects introduced by a structured environment. A natural extension is to replace the first–order time derivative in the Lindblad equation with a fractional derivative, producing a non–Markovian but still completely positive and trace–preserving (CPTP) evolution.
We consider the stationary fractional equation of motion
$\prescript{C}{}{D}_{t}^{\alpha}\rho(t)=\lambda\,\mathcal{L}\rho(t),\qquad 0<\alpha<1,$ (37)
where $\mathcal{L}$ is the standard Lindbladian generator. Note that Eq. (37) provides the formal CPTP generator: the fractional evolution is a subordination of the usual semigroup $e^{u\mathcal{L}}$ over an inverse–stable operational–time distribution Mainardi2000; Tarasov2021. For pure dephasing, $\mathcal{L}$ acts diagonally in the energy basis, and the off–diagonal element $u(t)=\rho_{01}(t)$ satisfies
$\prescript{C}{}{D}_{t}^{\alpha}u(t)=-\lambda u(t),\qquad u(0)=1.$ (38)
Taking Laplace transforms on both sides gives
$s^{\alpha}\tilde{u}(s)-s^{\alpha-1}=-\lambda\tilde{u}(s)~~\Rightarrow~~\tilde{u}(s)=s^{\alpha-1}/(s^{\alpha}+\lambda),$ (39)
whose inverse transform yields the Mittag–Leffler relaxation,
$u_{\alpha}(t)=E_{\alpha}(-\lambda t^{\alpha}),$ (40)
where $(\alpha,\lambda)$ is a scalar pair that can be optimized for reproducing the expected dephasing. The same Mittag–Leffler envelope has recently appeared in the context of the generalized fractional Rabi problem, where fractional differentiation modulates the coherent oscillations of a driven two-level system lopez2025generalisedfractionalrabiproblem.
Figure 4 compares the coherence amplitude $|\langle\sigma_{+}(t)\rangle|$ obtained from the single–order fractional propagation (40) with the numerically exact NZ evolution for representative bath exponents $\chi=0.5$, $1.0$, and $1.5$. For each case, the fractional parameters $(\alpha,\lambda)$ are optimized within the shaded intermediate–time windows and then the parameters are fixed for the fractional propagation over the entire time domain. Concretely, $(\alpha,\lambda)$ are obtained by minimizing the discrete root–mean–square deviation between the fractional coherence $u_{\alpha}(t;\alpha,\lambda)$ and the NZ result $u_{\mathrm{NZ}}(t)$ over a fitting interval $t_{\mathrm{start}},t_{\mathrm{end}}]$, cf. Appendix D. The fitting windows are chosen to be several times the bath correlation time $\tau_{B}\sim 1/\omega_{c}$ (so that both the initial curvature and the onset of the long–time tail are resolved) yet shorter than the recurrence time set by the high–frequency cutoff.
This consistent fitting procedure is applied to the sub–Ohmic, Ohmic, and super–Ohmic examples shown in the top, middle, and bottom panels, respectively. As can be seen, the fractional model accurately reproduces the exact coherence for the sub–Ohmic bath (with the deviation <0.02<0.02 over the entire time domain regardless of the selected time windows for the parameter fitting), demonstrating that a single fractional order can faithfully mimic the memory kernel encoded in $Q(t)$. Since $Q(t)$ is completely determined by the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ and hence by the spectral density $J(\omega)$, the optimized $(\alpha,\lambda)$ can be viewed as compact, two–parameter summaries of the influence of $C(t)$ over the chosen time window. For the Ohmic and super–Ohmic baths, the agreement gradually worsens as $\chi$ increases, reflecting the reduced influence of long–tailed correlations. On the other hand, the accuracy in the Ohmic baths can be systematically improved by fitting over a broader time window, which allows the optimized $(\alpha,\lambda)$ to capture both the short–time curvature and the slower asymptotic decay of the exact solution. Nevertheless, improvements for the super–Ohmic bath are modest compared to the other two baths, and its large residual deviation (>0.05>0.05) motivates the modified fractional ansatz, which incorporates the finite plateau $u_{\infty}$ characteristic of super-Ohmic environments.
For super–Ohmic baths ($\chi>1$), the coherence does not vanish at long times but saturates to a finite plateau $u_{\infty}=e^{-Q_{\infty}}$ determined by the bath reorganization energy. To preserve this limit, the fractional model must include $u_{\infty}$ explicitly through a convex mixture of the Markovian and fractional channels,
$u(t)=u_{\infty}+(1-u_{\infty})\,E_{\alpha}(-\lambda t^{\alpha})\quad\Rightarrow\quad v(t)=\frac{u(t)-u_{\infty}}{1-u_{\infty}}=E_{\alpha}(-\lambda t^{\alpha}).$ (41)
This affine form bounds the coherence between $u_{\infty}$ and 1 and ensures the correct long–time plateau. We therefore fit the plateau–normalized curve $v(t)$ within the shaded intermediate–time window to obtain $(\alpha,\lambda)$, and then reconstruct $u(t)$ from Eq. (41). The optimized fractional orders are systematically larger than s1s-1, compensating for the overly rapid decay of the fixed–order model and capturing the slower relaxation characteristic of structured baths. As shown in Figure 5, this plateau–anchored fractional ansatz accurately reproduces both the intermediate–time curvature and the long–time coherence plateau for all super–Ohmic exponents considered ($\chi=1.2$, $1.5$, and $1.8$).
Figure 5: Improved single–order fractional propagation for super–Ohmic baths with spectral exponents $\chi=1.2$, $1.5$, and $1.8$. Fractional propagations are anchored to the finite coherence plateau $u_{\infty}$ (see Eq. (41)). The fractional orders $(\alpha,\lambda)$ are obtained by least–squares fitting to the numerically exact coherence (solid lines, computed by direct integration of Eq. (26)) within the shaded intermediate–time window $t\!\in\![2/\omega_{c},20/\omega_{c}]$, and then the parameters are fixed for the fractional propagation over the entire time domain. Circle markers denote the fractional results, while the right panel shows the absolute deviation from the numerically exact result (solid lines, computed by direct integration of Eq. (26)).
Although $(\alpha,\lambda)$ can be determined by least–squares fitting, a simple physics–based choice can avoid nonlinear optimization entirely. From the sub–Ohmic kernel, the Laplace-domain relation $\mathcal{L}\{t^{-\beta}\}\!\sim\!s^{\beta-1}$ implies that a convolution kernel $t^{-\beta}$ entering the memory term $\int_{0}^{t}\!(t-\tau)^{-\beta}u(\tau)\,d\tau$ is equivalent to a fractional derivative of order $1-\beta$ Podlubny1999; Kilbas2006. Hence, the asymptotic behavior $Q^{\prime}(t)\!\sim\!t^{-\chi}$ yields the scaling rule $\alpha\simeq 1-\chi$, which ensures that a fractional convolution reproduces the long–memory tail of the exact $Q(t)$. In practice, when the accessible simulation window lies in the pre–asymptotic regime, the best effective order is slightly larger ($\alpha_{\mathrm{eff}}\!\in(1-\chi,1)$), capturing the stretched–exponential envelope of $u_{\mathrm{NZ}}(t)$ before the Mittag–Leffler power law sets in. A quick, data–driven estimate can be obtained from the local log–log slope of $X(t)=-\ln|u_{\mathrm{NZ}}(t)|$ (for sub-Ohmic and Ohmic baths) or $X(t)=-\ln|v(t)|$ (for super-Ohmic bath) as $\alpha_{\text{loc}}(t_{i})=\big[\ln X(t_{i+1})-\ln X(t_{i-1})\big]/\big[\ln t_{i+1}-\ln t_{i-1}\big],$ with $\alpha$ chosen as the median over the region where the slope is approximately constant. Once $\alpha$ is fixed, $\lambda$ follows from a single point in the same window by solving Eq. (40) or (41). This optimization–free selection rule recovers the expected scaling $\alpha\!\approx\!1-\chi$ for sub–Ohmic baths and provides reliable initial guesses for the least–squares refinement used in Figs. 4 and 5.
It is worth noting that the single–order fractional Lindbladian in Eq. (37) serves as a minimal demonstration because it admits exact analytical benchmarking against the solvable pure–dephasing model. Nevertheless, the underlying construction is fully general: the subordination representation (17) applies to any bounded generator $\mathcal{L}$, including dissipative channels that induce population decay, and to composite or multi–qubit systems via tensor–product generators $\mathcal{L}_{\text{tot}}=\sum_{i}\mathcal{L}_{i}\otimes\mathbb{I}_{\bar{i}}$. In these cases, the fractional propagation remains CPTP whenever each $\mathcal{L}_{i}$ is of GKSL form. The subordination framework thus extends beyond pure dephasing and accommodates arbitrary combinations of coherent and incoherent processes. A systematic generalization to distributed–order forms $\int_{0}^{1}\prescript{C}{}{D}_{t}^{\alpha}\rho\,g(\alpha)\,d\alpha=\mathcal{L}\rho$ is straightforward and provides a flexible route to model strongly structured environments.
IV.3 Unified Interpretation and Quantum Algorithms for Fractional Dynamics
The fractional framework is able to provide both a unified interpretation of non-Markovian quantum dynamics and a constructive pathway toward their quantum simulation. Numerical benchmarks on the spin–boson model show that a stationary single-order fractional Lindbladian can closely track the exact Nakajima–Zwanzig coherence over broad time windows, with errors at the few-percent level. Operationally, the fractional order $\alpha$ continuously tunes the degree of memory: lowering $\alpha$ slows relaxation, broadens correlation times, and induces algebraic envelopes, all consistent with the long-tail structure inherited from the bath spectral density $J(\omega)$. Within this hierarchy, the Liouville, Lindblad, and fractional master equations correspond respectively to unitary, Markovian, and structured non-Markovian evolution. Fractional differentiation introduces memory directly at the generator level, embedding power-law temporal correlations without violating complete positivity, and providing a continuous interpolation between reversible ($\alpha\!=\!1$) and strongly non-Markovian ($\alpha\!\ll\!1$) limits.
The subordination representation not only clarifies the physical origin of long-memory relaxation but also provides a direct blueprint for quantum simulation. In particular, the Bochner–Phillips subordination representation (17) replaces explicit memory integration by an average over standard Lindblad semigroups weighted by the inverse-stable density $f_{\alpha}(u,t)$. This representation guarantees complete positivity and naturally inspires two complementary quantum-simulation paradigms.
(i) Polynomial approximation via QSP/QSVT.
Since the Mittag–Leffler function $E_{\alpha}(z)$ is entire, the fractional propagator $E_{\alpha}(\mathcal{L}t^{\alpha})$ can be uniformly approximated by bounded polynomials on $\mathrm{spec}(\mathcal{L})$ and implemented through Quantum Signal Processing (QSP) or Quantum Singular Value Transformation (QSVT) LowChuang2019; Gilyen2019; Childs2018; Liu2021, generalizing exponential simulation.
Let $P_{N}(z)=\sum_{n=0}^{N}c_{n}z^{n}$ with $c_{n}=1/\Gamma(\alpha n+1)$. For a block-encoded generator
$U_{\mathcal{L}}=\begin{pmatrix}\mathcal{L}/\lambda_{\max}&\cdot\\[3.0pt] \cdot&\cdot\end{pmatrix}$ (42)
(with spectral radius $\lambda_{\max}$) acting on an extended Hilbert space such that $\langle 0|U_{\mathcal{L}}|0\rangle=L/\lambda_{\max}$, $P_{N}(\mathcal{L}t^{\alpha}/\lambda_{\max})$ realizes an approximate Mittag–Leffler evolution with deterministic error
$\|E_{\alpha}(z)-P_{N}(z)\|\leq|z|^{N+1}/\Gamma[\alpha(N+1)+1].$ (43)
A degree $N=\mathcal{O}(\sqrt{t^{\alpha}\lambda_{\max}\log(1/\epsilon)})$ achieves precision $\epsilon$, giving polylogarithmic scaling in $1/\epsilon$. This route extends exponential-type algorithms to non-Markovian propagators while retaining the CPTP structure of the underlying Lindblad generator.
(ii) Subordination sampling (fractional quantum trajectories).
Alternatively, one may sample operational times $u$ from $f_{\alpha}(u,t)$ and simulate standard Lindblad evolutions $e^{u\mathcal{L}}\rho(0)$ for each trajectory. Every path corresponds to a physical CPTP evolution of random duration, and observable estimates
$\langle O\rangle_{t}=\frac{1}{M}\sum_{k=1}^{M}\mathrm{Tr}[O\,e^{u_{k}\mathcal{L}}\rho(0)]$ (44)
converge as $\mathcal{O}(M^{-1/2})$ with unbiased variance. This defines a fractional quantum-trajectory method that emulates long-memory dynamics using only Markovian primitives Suess2014 and requires only constant memory per trajectory, since no time-history storage is needed.
In terms of complexity and accuracy, classical fractional solvers scale quadratically in propagation length because they store full history. Both quantum strategies circumvent this cost: the QSP/QSVT approach yields deterministic super-exponential error decay in $N$ and polylogarithmic gate complexity, whereas the subordination-sampling method scales linearly in trajectory number $M$ (with $M=\mathcal{O}(1/\epsilon^{2})$ trajectories sufficient to estimate bounded observables to accuracy $\epsilon$, essentially independent of system size) and polylogarithmically in Lindblad simulation cost. Their complementary error behaviors, that is deterministic for QSP and stochastic $1/\!\sqrt{M}$ for sampling, make them respectively suited for fault-tolerant and near-term architectures. We note that, in principle, the Lindblad evolution $e^{u\mathcal{L}}$ could also be implemented approximately via randomized circuits and engineered noise channels dambal2025harnessingintrinsicnoisequantum, suggesting an interesting direction for combining fractional quantum trajectories with noise engineering on NISQ hardware.
These approaches demonstrate that long-memory quantum processes can be simulated without explicit history storage. By either polynomially approximating the Mittag–Leffler propagator or stochastically averaging over Lindblad evolutions, fractional dynamics emerge as physically consistent CPTP mixtures of Markovian semigroups. This dual theoretical–algorithmic perspective connects open system theory with practical quantum-simulation paradigms, laying the groundwork for scalable studies of non-Markovian dynamics, dissipation, and decoherence within existing hardware toolkits.
While the fault-tolerant strategies discussed above—polynomial approximation via QSP/QSVT and stochastic subordination sampling—provide provably accurate and asymptotically efficient realizations of fractional propagation, recent studies have also pursued near-term hybrid alternatives. In particular, Leong et al. Leong2024 introduced a variational quantum–classical algorithm for solving Caputo time-fractional differential equations, employing overlap history states to capture memory effects on noisy hardware. Such variational schemes are complementary to the present framework: they enable proof-of-concept demonstrations of fractional evolution on NISQ devices, whereas the block-encoded and sampling-based approaches developed here are designed for scalable, fault-tolerant architectures with rigorous accuracy and resource guarantees.
V Conclusion and Outlook
Fractional calculus provides a rigorous and physically consistent extension of quantum master equations for simulating non-Markovian dynamics on both classical and quantum hardware. It embeds long-memory effects within a completely positive framework. By deriving the fractional master equation, proving its equivalence to power-law convolution forms, and expressing it through Lindblad subordination, we have unified three cornerstones of open-system theory—Liouville, Lindblad, and structured non-Markovian evolution—within a single operator-theoretic hierarchy.
Beyond the single-system formulation, the probabilistic subordination picture naturally extends to composite generators and multiqubit systems, since the subordination integral in Eq. (17) acts operatorwise and preserves tensor–product structure. Hence, the assumption of a GKSL generator is a sufficient but not restrictive condition for guaranteeing complete positivity; non-GKSL extensions can also be analyzed provided the underlying semigroup $e^{u\mathcal{L}}$ remains positive and norm-bounded.
It is also instructive to contrast this subordination hierarchy with the well-known recursion hierarchies of memory-kernel and time-convolutionless (TCL) master equations Breuer2016; Chruscinski2017; deVega2017. In traditional formulations, higher-order kernels arise from perturbative projections of the system–bath Liouvillian, and complete positivity is generally not preserved at finite truncation order. By contrast, the present construction defines a stochastic-process hierarchy in which each level corresponds to a CPTP map generated by a Bochner–Phillips subordination of a Lindblad semigroup. The resulting hierarchy organizes non-Markovian dynamics in terms of nested waiting-time distributions rather than nested kernels, providing a physically transparent and CPTP-preserving alternative to conventional kernel recursion schemes.
This unified picture clarifies how fractional differentiation introduces structured memory directly at the generator level, producing CPTP yet non-divisible dynamics that interpolate smoothly between unitary, Markovian, and strongly non-Markovian regimes. It connects the algebraic tails of long-time relaxation with the complete monotonicity of the subordination kernel and explains why fractional dynamics form a structured subclass of memory-kernel quantum master equations.
From an algorithmic perspective, the same structure offers a practical route to simulation. The Mittag–Leffler propagator can be approximated deterministically via polynomial quantum signal processing or stochastically through subordination sampling over operational times. Both approaches realize long-memory quantum dynamics using only Markovian primitives, avoiding explicit time-history storage while preserving complete positivity. This dual deterministic–stochastic formulation establishes a bridge between theoretical open-system analysis and scalable quantum algorithms.
Although the present analysis focused on the exactly solvable pure-dephasing limit at zero temperature, the fractional framework itself imposes no such restriction. Because the subordination construction remains valid for any GKSL generator, the same approach can incorporate amplitude damping, population relaxation, and multiqubit tensor-product generators while preserving complete positivity. Preliminary tests also indicate that the fitted fractional parameters exhibit smooth temperature dependence, suggesting that a single fractional order can continue to describe long-memory behavior for $k_{B}T/\omega_{c}\lesssim 1$.
Future research will extend these ideas to distributed-order and space–time-fractional models that capture correlated noise and anomalous diffusion, derive rigorous error bounds for quantum implementations of Mittag–Leffler propagators, and develop hybrid quantum–classical workflows that exploit fractional subordination for efficient noise modeling and dissipation control. More broadly, the fractional framework provides a co-designed theoretical language linking operator theory, stochastic processes, and quantum simulation—offering a unified platform for exploring memory, coherence, and irreversibility in complex quantum systems.
Acknowledgements.
This work is supported by a U.S. Department of Energy, Office of Science, Early Career Research Program award in the Basic Energy Sciences, Division of Chemical Sciences, Geosciences, and Biosciences, Computational and Theoretical Chemistry program under FWP 83466. Pacific Northwest National Laboratory (PNNL) is operated by Battelle for the U.S. DOE under contract DE-AC05-76RL01830. Y.Z. acknowledges the support from Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (contract no. 89233218CNA000001).
Data Availability Statement
The data that support the findings of this study are available within the article [and its supplementary material].
Appendix A Caputo–Volterra Equivalence and Convolution-Type QME
We collect the basic operators and prove the equivalence between the Caputo fractional master equation and its Volterra (integral) and convolution (differential) forms, keeping careful track of boundary terms. The equivalence between the Caputo, Volterra, and convolution forms follows standard results in fractional calculus Podlubny1999; Kilbas2006; Luchko2020 and their applications to open quantum systems Tarasov2021. The complete monotonicity of the Volterra kernel is a direct consequence of Bernstein’s theorem and the classical subordination results of Bochner and Feller Bochner1955; Feller1971.
A.1 Operators and identities
For $\mu>0$, the Riemann–Liouville (RL) fractional integral is given by
$(I_{t}^{\mu}f)(t)=\frac{1}{\Gamma(\mu)}\int_{0}^{t}(t-\tau)^{\mu-1}f(\tau)~d\tau,$ (45)
and for $0<\alpha<1$, the corresponding Caputo derivative is
$\big(\prescript{C}{}{D}_{t}^{\alpha}f\big)(t)$ $=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}(t-\tau)^{-\alpha}f^{\prime}(\tau)~d\tau$
$=I_{t}^{1-\alpha}f^{\prime}(t).$ (46)
They satisfy, for sufficiently regular function $f$,
$I_{t}^{\mu}I_{t}^{\nu}$ $=I_{t}^{\mu+\nu},$
$\prescript{C}{}{D}_{t}^{\alpha}\circ I_{t}^{\alpha}f$ $=f,$ (47)
$I_{t}^{\alpha}\circ\prescript{C}{}{D}_{t}^{\alpha}f$ $=f-f(0).$
A.2 Equivalence of forms for the fractional Lindblad equation
Proposition 1 (Caputo–Volterra–Convolution equivalence).
Let $0<\alpha<1$, $\rho$ be absolutely continuous, and $\mathcal{L}$ a (bounded) GKSL generator on the working space. The following are equivalent:
  1. 1.
    (Caputo):
    $\prescript{C}{}{D}_{t}^{\alpha}\rho(t)=\mathcal{L}\rho(t).$ (48)
  2. 2.
    (Volterra):
    $\rho(t)=\rho(0)+\int_{0}^{t}K_{\alpha}^{(\mathrm{V})}(t-\tau)~\mathcal{L}\rho(\tau)~d\tau,$ (49)
    with $K_{\alpha}^{(\mathrm{V})}(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ for $t>0$.
  3. 3.
    (Convolution, differential):
    $\dot{\rho}(t)=\int_{0}^{t}k_{\alpha}(t-\tau)~\mathcal{L}\rho(\tau)~d\tau+\frac{t^{\alpha-1}}{\Gamma(\alpha)}~\mathcal{L}\rho(0),$ (50)
    with $k_{\alpha}(t)=\frac{d}{dt}K_{\alpha}^{(\mathrm{V})}(t)=\frac{t^{\alpha-2}}{\Gamma(\alpha-1)}$ for $t>0$.
If $\mathcal{L}\rho(0)=0$ (e.g., $\rho(0)$ is a fixed point of $\mathcal{L}$), the boundary term vanishes and the differential form is a pure convolution.
Proof. (Caputo $\Rightarrow$ Volterra): Apply $I_{t}^{\alpha}$ to the Caputo equation and use (47): $\rho(t)-\rho(0)=I_{t}^{\alpha}[\mathcal{L}\rho](t)=\int_{0}^{t}K_{\alpha}^{(\mathrm{V})}(t-\tau)~\mathcal{L}\rho(\tau)~d\tau.$ (Volterra $\Rightarrow$ Convolution): Differentiate the Volterra integral (Leibniz rule for weakly singular kernels) to obtain $\dot{\rho}(t)=(k_{\alpha}*\mathcal{L}\rho)(t)+K_{\alpha}^{(\mathrm{V})}(t)~\mathcal{L}\rho(0).$ (Convolution $\Rightarrow$ Caputo): Integrate the differential form with $I_{t}^{1-\alpha}$, use $I_{t}^{1-\alpha}\dot{\rho}=\prescript{C}{}{D}_{t}^{\alpha}\rho$, and $I_{t}^{1-\alpha}k_{\alpha}=\delta$ in the distributional sense (since $k_{\alpha}=\frac{d}{dt}K_{\alpha}^{(\mathrm{V})}$), yielding $\prescript{C}{}{D}_{t}^{\alpha}\rho=\mathcal{L}\rho$. ∎
Remark 2 (Three kernels and their roles).
There are three related kernels:
  • $\kappa_{C}(t)=\frac{t^{-\alpha}}{\Gamma(1-\alpha)},$
  • $K_{\alpha}^{(\mathrm{V})}(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)},$
  • $k_{\alpha}(t)=\frac{t^{\alpha-2}}{\Gamma(\alpha-1)}.$
Only $K_{\alpha}^{(\mathrm{V})}$ and $k_{\alpha}=\frac{d}{dt}K_{\alpha}^{(\mathrm{V})}$ appear as the memory kernels multiplying $\mathcal{L}\rho$ in the integral and differential QME forms, respectively; $\kappa_{C}$ multiplies $\dot{\rho}$ inside the Caputo derivative. Moreover, $K_{\alpha}^{(\mathrm{V})}$ is completely monotone on $\mathbb{R}_{+}$, hence is the Laplace transform of a positive measure; this underpins the subordination representation of the fractional semigroup.
The Caputo derivative used throughout this work is preferable for physical initial–value problems because it acts on the first derivative of $\rho(t)$, ensuring that the initial condition $\rho(0)$ enters linearly and that $\prescript{C}{}{D}_{t}^{\alpha}\rho(t)\!\to\!\dot{\rho}(t)$ as $\alpha\!\to\!1$. By contrast, the Riemann–Liouville derivative introduces fractional integrals of $\rho(t)$ that depend explicitly on its entire prehistory, requiring fractional initial conditions that are not directly measurable. In open system dynamics where the state $\rho(0)$ is known, the Caputo definition yields physically meaningful time evolution and consistent reduction to standard Lindblad dynamics in the limit $\alpha\!\to\!1$. Numerically, both definitions produce similar late–time algebraic behavior, but only the Caputo form preserves the conventional initial–state specification used in quantum master equations.
Appendix B Fractional Adams–Moulton (Volterra) Discretization in Banach Spaces
The numerical scheme presented here follows the fractional Adams–Moulton predictor–corrector formulation for Caputo equations Diethelm2002; Podlubny1999; Kilbas2006, widely used for fractional differential equations.
Specifically, we derive an implicit predictor–corrector (Diethelm–Ford–Freed type) scheme for
$\prescript{C}{}{D}_{t}^{\alpha}\rho(t)=\mathcal{L}\rho(t),\qquad 0<\alpha<1,$ (51)
with $\rho:[0,T]\to\mathcal{B}_{1}(\mathcal{H})$ absolutely continuous and $\mathcal{L}$ the GKSL generator of a norm-continuous CPTP semigroup. By Prop. 1 (Caputo $\Leftrightarrow$ Volterra),
$\rho(t)=\rho(0)+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-\tau)^{\alpha-1}~\mathcal{L}\rho(\tau)~d\tau.$ (52)
Let $t_{n}=nh$ with $h=T/N$, and denote $\rho_{n}\approx\rho(t_{n})$. Using piecewise linear (predictor) and trapezoidal (corrector) quadrature of (52) yields the standard Lubich weights:
$b_{j}=(j+1)^{1-\alpha}-j^{1-\alpha}\quad(j\geq 0),$ (53)
and composite corrector weights $a_{n-k}$ obtained by integrating exactly against $(t_{n+1}-\tau)^{\alpha-1}$ on each subinterval $t_{k},t_{k+1}]$:
$a_{n-k}=(n-k+1)^{\alpha}-2(n-k)^{\alpha}+(n-k-1)^{\alpha}$ (54)
with $1\leq k\leq n-1$, $a_{0}=1^{\alpha}-0^{\alpha}$, and $a_{n}=1^{\alpha}$ (endpoint corrections are the usual ones for the fractional trapezoid rule).
Employing the predictor–corrector update, the predictor is
$\rho_{n+1}^{\mathrm{(pred)}}=\rho_{0}+\frac{h^{\alpha}}{\Gamma(1+\alpha)}\sum_{k=0}^{n}b_{n-k}~\mathcal{L}\rho_{k}~,$ (55)
and the corrector is implicit:
$\rho_{n+1}=\rho_{0}+\frac{h^{\alpha}}{\Gamma(1+\alpha)}\left[\sum_{k=0}^{n}a_{n-k}~\mathcal{L}\rho_{k}+a_{-1}~\mathcal{L}\rho_{n+1}^{\mathrm{(pred)}}\right]$ (56)
with $a_{-1}=1$. Equivalently,
$\left(\mathbb{I}-\frac{h^{\alpha}}{\Gamma(1+\alpha)}~\mathcal{L}\right)\rho_{n+1}=\rho_{0}+\frac{h^{\alpha}}{\Gamma(1+\alpha)}\sum_{k=0}^{n}a_{n-k}~\mathcal{L}\rho_{k}.$ (57)
In a finite representation, the left factor is step-invariant and can be factorized once (e.g., using LU/Cholesky factorization).
[…]
The naive history cost per step is $\mathcal{O}(n)$. Efficient long-time propagation can be achieved using compressed-memory implementations based on sum-of-exponentials approximations Garrappa2015,
$\frac{t^{\alpha-1}}{\Gamma(\alpha)}\approx\sum_{q=1}^{Q}w_{q}e^{-\xi_{q}t},$ (58)
yields $\mathcal{O}(1)$ history updates via $Q$ auxiliary ODEs, for overall $\mathcal{O}(QN)$ time and $\mathcal{O}(Q)$ memory with controllable error.
Appendix C Subordination for the Fractional Master Equation
The subordination representation used in Sec. III.3 follows the classical theory of subordinated semigroups and inverse-stable processes Bochner1955; Feller1971; BaeumerMeerschaert2001; MeerschaertSikorskii2012.
We assume $\mathcal{L}$ is the (time-independent) GKSL generator of a norm-continuous CPTP semigroup $T(u)=e^{u\mathcal{L}}$ on the trace-class $\mathcal{B}_{1}(\mathcal{H})$ with growth bound $\omega_{0}\leq 0$. $\rho:[0,\infty)\to\mathcal{B}_{1}(\mathcal{H})$ is absolutely continuous and of exponential order. For $0<\alpha<1$, the Laplace transform of the Caputo derivative satisfies
$\mathbb{L}\big\{\prescript{C}{}{D}_{t}^{\alpha}\rho(t)\big\}(s)=s^{\alpha}\hat{\rho}(s)-s^{\alpha-1}\rho(0),$ (59)
where $\hat{\rho}(s)$ is the Laplace tramsform of $\rho(t)$
[…]
Applying this to Eq. (6) and using $\mathbb{L}\{\mathcal{L}\rho\}(s)=\mathcal{L}\hat{\rho}(s)$ (time-independence of the Lindblad operator) gives
$\big(s^{\alpha}\mathbb{I}-\mathcal{L}\big)\hat{\rho}(s)=s^{\alpha-1}\rho(0)$
$\Rightarrow~$ $\hat{\rho}(s)=s^{\alpha-1}\big(s^{\alpha}\mathbb{I}-\mathcal{L}\big)^{-1}\rho(0).$ (61)
[…]
$(\lambda\mathbb{I}-\mathcal{L})^{-1}=\int_{0}^{\infty}e^{-\lambda u}~T(u)~du.$ (62)
[…]
$\big(s^{\alpha}\mathbb{I}-\mathcal{L}\big)^{-1}=\int_{0}^{\infty}e^{-us^{\alpha}}~T(u)~du.$ (63)
Substituting this expression into Eq. (C) yields
$\hat{\rho}(s)=\int_{0}^{\infty}\Big[s^{\alpha-1}e^{-us^{\alpha}}\Big]~T(u)\rho(0)~du.$ (64)
The prefactor $s^{\alpha-1}e^{-us^{\alpha}}$ acts as the Laplace-domain representation of the kernel that connects the physical time $t$ and the operational time $u$. We now introduce this kernel explicitly by defining $f_{\alpha}(u,t)$ as the function whose Laplace transform in $t$ satisfies
[…]
This property follows from Bernstein’s theorem on completely monotone functions Schilling2012Bernstein. The function $f_{\alpha}(u,t)$ is known as the inverse-stable (or hitting-time) density associated with an $\alpha$-stable subordinator. It provides a probabilistic bridge between the deterministic operational time $u$ and the physical time $t$, capturing the long-memory and non-Markovian effects introduced by fractional differentiation. This kernel possesses several key properties that follow directly from Eq. (65):
  1. 1.
    Positivity: $f_{\alpha}(u,t)\geq 0$ for all $u,t\geq 0$, as $s^{\alpha-1}e^{-us^{\alpha}}$ is a completely monotone function of $s$.
  2. 2.
    Normalization: $\int_{0}^{\infty}f_{\alpha}(u,t)~du=1$ for each $t\geq 0$. Indeed,
    $\mathbb{L}~\left\{\int_{0}^{\infty}f_{\alpha}(u,t)~du\right\}(s)$ $=\int_{0}^{\infty}s^{\alpha-1}e^{-us^{\alpha}}~du$
    $=\frac{s^{\alpha-1}}{s^{\alpha}}$
    $=\frac{1}{s},$ (66)
    and the inverse Laplace transform of $1/s$ is unity.
Finally, the fractional evolution can be written in the subordination form
$\rho_{\alpha}(t)=\int_{0}^{\infty}f_{\alpha}(u,t)~T(u)\rho(0)~du,$ (67)
which expresses the fractional dynamics as a superposition of standard semigroup evolutions $T(u)\rho(0)$ weighted by the random-time density $f_{\alpha}(u,t)$. This representation provides both a rigorous analytical foundation and a clear probabilistic interpretation for fractional master equations. In particular, taking the inverse Laplace transform of (64) in $t$ (and using Fubini/Tonelli to exchange the Bochner and scalar integrals, justified by positivity and exponential bounds) yields Eq. (17), which is the desired subordination representation, representing a convex mixture of Lindblad semigroups and remains CPTP, in agreement with the criteria for completely positive dynamical maps discussed in Hall2014; Rivas2014. ∎
Appendix D Optimization of fractional parameters
For each bath exponent $\chi$, the single–order fractional model is specified by the propagator
$u_{\alpha}(t;\alpha,\lambda)=E_{\alpha}\!\big(-\lambda t^{\alpha}\big),$ (68)
which is compared to the numerically exact Nakajima–Zwanzig coherence $u_{\mathrm{NZ}}(t)=\exp[-Q(t)]$ obtained from the structured spectral density $J(\omega)$. The parameters $(\alpha,\lambda)$ are determined by a least–squares fit over a finite time window $t_{\mathrm{start}},t_{\mathrm{end}}]$:
$\mathrm{RMSE}^{2}(\alpha,\lambda)=\frac{1}{N_{\mathrm{fit}}}\sum_{k:\,t_{k}\in[t_{\mathrm{start}},t_{\mathrm{end}}]}\Big|u_{\alpha}(t_{k};\alpha,\lambda)-u_{\mathrm{NZ}}(t_{k})\Big|^{2},$ (69)
with $(\alpha,\lambda)$ chosen to minimize $\mathrm{RMSE}^{2}(\alpha,\lambda)$. In practice we use a coarse grid search followed by a local refinement (Nelder–Mead) to identify the optimal parameters.
The fitting window is guided by the bath correlation time. In our dimensionless units we set $\omega_{c}=1$, so the bath correlation time is of order $\tau_{B}\sim 1/\omega_{c}=1$. More quantitatively, we estimate $\tau_{B}$ from the decay of the bath correlation function $C(t)=\langle B(t)B(0)\rangle$ (or equivalently from $Q^{\prime\prime}(t)$) as the time at which $|C(t)|$ has dropped below $e^{-1}|C(0)|$. The fitting windows are then chosen as $t_{\mathrm{start}}\approx 2\tau_{B}$ and $t_{\mathrm{end}}\in[20\tau_{B},60\tau_{B}]$, which in practice corresponds to $2/\omega_{c},20/\omega_{c}][$$2/\omega_{c},60/\omega_{c}]$ in Fig. 4. This ensures that the optimized $(\alpha,\lambda)$ are extracted from a regime where the bath has largely decorrelated but before finite-cutoff recurrences become important.
Since $Q(t)$ is obtained by a double time integral over the bath correlation function $C(t)$, the optimized pair $(\alpha,\lambda)$ should be regarded as an effective two–parameter representation of the influence of $C(t)$ over the chosen fitting window. In this sense, $\alpha$ controls the effective long–time decay law of the coherence, while $\lambda$ sets the overall decoherence timescale. For general structured spectral densities there is no closed-form analytic relation between $C(t)$ and $(\alpha,\lambda)$; the fractional parameters summarize the memory characteristics encoded in $Q(t)$ rather than reproduce them exactly.